# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

### v1

@@ 2.2 Guiding SCoBots
The inspectable nature of SCoBots not only brings the benefit of improved human-understandability, but importantly allows for targeted human-machine interactions. In the following we describe two guidance approaches for interacting with the latent representations of SCoBots: concept pruning and object-centric rewarding.
Concept pruning. The type and amount of concepts that are required for a successful policy may vary greatly across tasks. For instance, object colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman. At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g. the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$). Moreover, for a single RL task, the need for specific concepts might even change during training, e.g. in progressive environments where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2021).
Specifically, SCoBot’s human-comprehensible concepts allow domain experts to prune unnecessary concepts. In this way expert users can guide the learning process towards relevant concepts. Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$. Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, and (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$. iSCoBots thus formally refine Eq. 2, leading to:
@@
Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections. Human-normalized scores (red line) of deep and object-centric agents trained using PPO on 99 ALE environments. SCoBots obtain similar or better performance than the deep agents, showing object-centric agents can also learn to solve RL tasks while making use of human-understandable concepts (left). Guiding SCoBots allow to correct misalignment in Pong (center) and the ill-defined reward signal of Kangaroo (right).
To give brief examples of these four pruning possibilities we refer to Fig. 2, focusing here on the orange subsections. Particularly, in Pong a user can remove objects (scores) or specific object properties (R,G,B values) which would result in a $\overline{\Omega_{t}}$. Second, the color relation, color($\cdot$), is not relevant for successfully playing Pong and can therefore be removed entirely from $\mathcal{F}$. Third, a user can state that the vertical distance function (dist($\cdot,\cdot$).y) should not be applied to the (player, enemy) input couple. These pruning actions result in the set of relational concepts $\overline{\Gamma_{t}}$. Lastly, the only playable actions in Pong are UP and DOWN. To ease the search, the available FIRE, LEFT and RIGHT from the base Atari action space might be discarded as they are equivalent to NOOP (i.e. no action), resulting in $\overline{\mathcal{A}}$.
Importantly, being able to prune concepts can help mitigate misalignement issues such as those of agents playing Pong (cf. Figure 1), where an agent can base it’s action selection on the enemy’s position. Specifically, by pruning the enemy position from the distance relation concept a user can guide the SCoBot to rely on relevant features for their decisions such as the ball’s position, rather than the spurious correlation with the enemy’s paddle.
Object-centric feedback reward. Reward shaping (Touzet, 1997; Ng et al., 1999) is a standard RL technique that is used to facilitate the agent’s learning process by introducing intermediate reward signals, aggregated to the original reward signal. The object-centric and importantly concept-based approach of SCoBots allows to easily craft additional reward signals. Formally, one can use the extracted relational concepts to express new expert reward signal:
@@
where $R^{\text{exp}}:\Gamma\longrightarrow\mathbb{R}$. Moreover, where $\Gamma_{t}=\mu(\omega(s_{t}))$ is the relational state, extracted by the relation extractor of a SCoBot. The coefficient $\alpha_{g_{t}}\in\mathbb{R}$ is used to penalize or reward the agent proportionally to relational concepts. Our expert reward only relies on the state, as we make use of the concepts extracted from it. However, incorporating the action into the reward computation is straightforward. In practice, this expert reward signal can lead to guiding the agent towards focusing on relevant concepts, but also help smoothing sparse reward signals, which we will discuss further in our evaluations.
@@
For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf. red components of Fig. 2), with the intention of incentivizing the agent to maintain close proximity with the ball. As we show in our experimental evaluation, we can use this concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf. App. A.9 for details on each problem).
3 Experimental Evalautions
@@
In our evaluations, we investigate several properties and potential benefits of SCoBots. We specifically aim to answer the following research questions:
@@
  • Q1. Are SCoBot agents able to learn competitive policies on benchmark RL environments?
@@
  • Q2. Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes, e.g. policy misalignments?
@@
  • Q3. Can concept-based guidance let us mitigate RL caveats of SCoBot agents, e.g. policy misalignments?
@@
Figure 4: SCoBots allow to follow their decision process. Their object centricity and interpretable concepts allow us to follow their reasoning. The decision processes of SCoBots (extracted from the decision trees) are provided with their associated states for an guided SCoBot on Skiing (left) and a SCoBot on Pong (right). For examble, in this Skiing state, our iSCoBot selects RIGHT, as the distance between Player and the (left) Flag1 (on the $x$ axis) is high (bigger than $15$). This agent selects the correct action for the right reasons.
@@
Experimental setup: We evaluate SCoBots on $9$ Atari games (v5 version) from the ALE benchmark (Bellemare et al., 2012) (by far the most used RL framework (cf. App. A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game. For results based on achieved game scores we provide human normalized scores (following eq. 5) that are averaged over $3$ seeds for each agent configuration. We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and specify these as neural agents in the following. Note that we evaluate these agents on the latest v5 version of the environments, leading to slightly worst results than (Schulman et al., 2017). We obtain similar results when evaluating on the old versions (cf. Tab. 1). For the SCoBots we assume access to a pretrained object extractor and thus provide our agents object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a). This allows us to focus our evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g. (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)). For all agents, we restrict training to $20\text{M}$ frames and use the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 (Raffin et al., 2021) implementation and its default hyperparameters for both the deep neural and SCoBot agents (cf. Tab. 2 in the Appendix). SCoBots’ inputs consist of vectors containing the values of concatenated concepts (i.e. properties and relations). Specifically, in our evaluations, the set of object properties consists of object class (e.g. enemy paddle, ball etc.), $(x^{t},y^{t})$ coordinates, coordinates at the previous position $(x^{t-1},y^{t-1})$, height and width, the most common RGB values (i.e. the most representative color of an object), and object orientation. The specific set of functions, $\mathcal{F}$, that are used to extract object relations is composed of: euclidean distance, directed distances (on $x$ and $y$ axes), speed, velocity, plan intersection, the center (of the shortest line between two objects), and color name (cf. App. A.6.1 for implementation details). To distill SCoBots learned policy of SCoBots, we use the rule-based model extraction method from (Zarlenga et al., 2021). For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf. Tab. 4). Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below. More details about the experimental setup and the hyperparameters are provided in App. A.6.
SCoBots learn competitive policies (Q1). We present human-normalized (HN) scores of both SCoBot and neural agents trained on each investigated game individually in Fig. 3 (left). Exact numerical results are provided in Table 1 (cf. App. A.3). Overall, our results suggest that SCoBots provide competitive performances on every tested game despite the constraints of multiple bottleneck within their architectures. This shows algorithms where the object extraction is decoupled from the reasoning process that leads to the action selection can be developed. Other environments might require additional information extraction processes (e.g. MsPacman, in which the agent must navigate a maze), and we haven’t evaluated on such environments here. This constitute a limit to our SCoBots.
@@
Inspectable SCoBots’ to detect misalignments (Q2). A main target of developing ScoBots is to obtain competitive, yet transparent RL agents that allow for human users to identify an agent’s underlying reasoning process. Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allow to pinpoint not just the object properties, but importantly also the relational concepts and properties that are being used for the final action decision. This is exemplified in Fig. 4 for the games of Skiing and Pong (cf. App. A.5 for explanations of SCoBots on the remaining games). Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of the two games. For example, for the specific game state in Skiing, the SCoBot agent selects the action RIGHT as the best option, because the distance from its character to the left flag is larger than a specific value ($15$ pixels). Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
@@
A more striking display of the benefits the inherent explanations of SCoBot agents can be found in the Pong-based agent in Fig 4. Specifically, the provided explanation suggests that the agent is basing its decision on the vertical position of the enemy’s and of its own paddle (distance between the two paddles y direction). This, in turn, suggests that in fact the agent is ignoring the ball’s position, at least to a certain degree. Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay). Thus, its vertical position is highly correlated with the enemy’s one (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf. App. A.8). Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.
@@
To validate these findings further we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (we refer to this environment as NoEnemy cf. Fig. 1). We evaluate both the neural and SCoBot agents that were trained on the original Pong environment and observe catastrophic performance drops in HN scores (cf. Fig. 3 both of SCoBot and both neural PPO) at a level of random policies, when evaluated on this environment (NoEnemy). We have also run an DQN agent on this environments and observed the same results. The drop in neural agent performance is particularly striking as initial importance maps produced by these neural agents had highlighted all $3$ moving objects as relevant for their decisions (i.e. the player, ball, and enemy) (cf. Fig.1 top left and App. A.8, aligned with findings of (Weitkamp et al., 2018)). These importance maps suggest the agents had based their decision on right reasons. Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and but more importantly the granularity of such post-hoc explainability methods in the context of RL. It highlights the importance of inherently explainable RL models on the level of concepts as provided via SCoBots to identify potential issues such as goal misalignment. In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
@@
We here make use of SCoBots interaction capabilities to address $4$ famous RL specific problems: goal misalignment which we had previously identified via SCoBot’s inherent, concept-based explanations and also sparse rewards, ill-defined rewards and difficult credit assignment.
@@
Realigning SCoBots: For mitigating identified SCoBots goal misalignment issues, we can correct the previously introduced misalignment, by simply removing (pruning) the enemy in the set of considered objects ($\overline{\Omega}$). Doing so, the enemy cannot be used for the action selection process. This leads to an agent that is able to play Pong and its NoEnemy version (guided SCoBot in Fig 3 (center)). This shows again that playing Pong without observing the enemy is achievable, and thus should also be the case for deep agents.
Sparse Reward RL: Let us move on to the issue of sparse reward in the context of RL. In Pong, to score a point, the player needs to return the ball and have it go past the enemy. To do so, the player has to perform a spiky shot, obtained by touching the ball on one of its paddle’s sides (otherwise, the shot is flat and the enemy is easily catching it). The probability of such a shot is approximately $3.8\%$ (as shown in details in Appendix A.2). Thus, a randomly playing (or untrained) agent is only rewarded on average every $2200$ steps. Providing a SCoBot agent with an additional reward that is inversely proportional to the distance between its paddle and the ball incentivizes the agent to keep a vertical position close to the ball’s one, tripling the probability of a winning shot in our experiment (cf. App. A.2). Thus, the reward signal in Pong is low, but our example illustrates how the interpretable concepts allow to easily guide SCoBots, making up for the reward sparsity.
@@
Ill-defined reward: As previously explained, defining a reward upfront that incentivizes agents to learn the expected behavior can be difficult. An example of ill-defined reward (i.e. a reward that will lead to an unaligned behavior if the agent maximizes the return) is present in Kangaroo. According to the documentation of the game111www.retrogames.cz/play_195-Atari2600.php, “the mother kangaroo on the bottom floor tries to reach the top floor where her joey is being held captive by some monkeys”. However, punching the monkey enemies gives a higher cumulative reward than climbing up to save the baby. RL agents thus tend to learn to kick monkeys on the bottom floor and never go to the joey (cf. Fig. 4). For revising such agents (i.e. incentivize them to reach the games’ original intended goal), we provide the SCoBot agents with an additional reward signal, based on the agent’s progression to the joey, (cf. App. A.9). As can be seen in Fig. 3 this reward allows us to obtain SCoBot agents that achieve high HN game scores Fig. 3 (left) that indeed complete the level Fig. 3 (right). In contrast the neural agents achieve high HN scores, but do not complete the levels, whereas unguided SCoBot agents achieve lower HN scores, but higher level completion ratios than the neural agents, but greatly lower ones than the guided SCoBot agents.
@@
Difficult Credit Assignment Problem: Lastly, the difficult credit assignment problem in RL is the challenge of correctly attributing credit or blame to past actions when the agent’s actions can have delayed and non-obvious effects on its overall performance. We illustrate this in the context of Skiing, where in standard configurations agents receive at each step a negative reward that corresponds to the number of seconds elapsed since the last steps (i.e. varying between 3-3 and 7-7). This reward aims at punishing agents for going down the slope slowly. Additionally, the game keeps track of the number of flags that the agent has correctly traversed (displayed at the top of the screen, cf. Fig. 4). A large positive reward, proportional to it, is attributed to the agent when it reaches the final state. Associating this large reward signal with the passed flags is extremely difficult, without prior knowledge on human skiing competitions. At each timestep, in the next evaluations we thus provide SCoBot agents with a reward that is proportional to the agent’s progression to the flags’ position to incentivize an agent to pass in between them as:
$R^{\text{exp}}_{1}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}-D(Player,o)^{t-1}.$
$R^{\text{exp}}_{2}=V(Player).y.$
@@
These rewards are then summed to the environmental one. With this we obtain guided SCoBot agents that reach human scores, whereas the neural and non-interactive SCoBot agents perform worse than randomly (cf. Fig. 3 (left)).
@@
Overall, our experimental evaluations show that the not only are SCoBots on par with neural RL agents, but that importantly the concept-based and inspectable nature of SCoBots allows identify and revise several important RL specific caveats. In the course of our evaluations we could identify several previously unknown issues of RL agents on benchmark environments even as simple as Pong.

### v2

@@ 2.2 Guiding SCoBots
The inspectable nature of SCoBots not only brings the benefit of improved human-understandability, but importantly allows for targeted human-machine interactions. In the following, we describe two guidance approaches for interacting with the latent representations of SCoBots: concept pruning and object-centric rewarding. We refer to the revised SCoBot agents as guided SCoBots in the following.
Concept pruning. The type and amount of concepts that are required for a successful policy may vary greatly across tasks. For instance, the objects colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman. At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g. the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$). Moreover, for a single RL task, the need for specific concepts might even change during training, e.g. in progressive environments (where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024)).
SCoBot’s human-comprehensible concepts therefore allow domain experts to prune unnecessary concepts. In this way, expert users can guide the learning process towards relevant concepts. Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$. Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, or (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$. Guided SCoBots thus formally refine Eq. 2, leading to:
@@
Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections. Human-normalized scores of deep and SCoBots (with and without guidance) agents trained using PPO on 99 ALE environments. SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left). Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level complition score of 100% on the intended goal’s evaluation in Kangaroo (right).
To give brief examples of these four pruning possibilities we refer to Fig. 2, focusing here on the blue subparts. Particularly, in Pong a user can remove objects (e.g. scores) or specific object properties (e.g. R,G,B values) to obtain $\overline{\Omega_{t}}$. Second, the color relation, color($\cdot$), is irrelevant for successfully playing Pong and can therefore be removed from $\mathcal{F}$. Third, the vertical distance function (dist($\cdot,\cdot$).y) can be prevented from being applied to the (player, enemy) input couple. These pruning actions provide SCoBots with $\overline{\Gamma_{t}}$. Lastly, the only playable actions in Pong are UP and DOWN. To ease the policy search, the available FIRE, LEFT and RIGHT from the base Atari action space might be discarded to obtain $\overline{\mathcal{A}}$, as they are equivalent to NOOP (i.e. no action).
Importantly, being able to prune concepts can help mitigate misalignement issues such as those of agents playing Pong (cf. Fig. 1), where an agent can base its action selection on the enemy’s position instead of the ball’s one. Specifically, pruning the enemy position from the distance relation concept enforces SCoBot agents to rely on relevant features for their decisions such as the ball’s position, rather than the spurious correlation with the enemy’s paddle.
Object-centric feedback reward. Reward shaping (Touzet, 1997; Ng et al., 1999) is a standard RL technique that is used to facilitate the agent’s learning process by introducing intermediate reward signals, aggregated to the original reward signal. The object-centric and importantly concept-based approach of SCoBots allows to easily craft additional reward signals. Formally, one can use the extracted relational concepts to express a new expert reward signal:
@@
where $R^{\text{exp}}:\Gamma\longrightarrow\mathbb{R}$ and $\Gamma_{t}=\mu(\omega(s_{t}))$ is the relational state, extracted by the relation extractor. The coefficient $\alpha_{g_{t}}\in\mathbb{R}$ is used to penalize or reward the agent proportionally to relational concepts. Our expert reward only relies on the state, as we make use of the concepts extracted from it. However, incorporating the action into the reward computation is straightforward. In practice, this expert reward signal can lead to guiding the agent towards focusing on relevant concepts, but also help smoothing sparse reward signals, which we will discuss further in our evaluations.
@@
For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf. “expert reward” arrow in Fig. 2), with the intention of incentivizing the agent to maintain close proximity with the ball. As we show in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf. App. A.8 for details on each problem).
3 Experimental Evaluations
@@
In our evaluations, we investigate several properties and potential benefits of the SCoBot agents. We specifically aim to answer the following research questions:
@@
(Q1) Are concept based agents able to learn competitive policies on different RL environments?
@@
(Q2) Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes?
@@
(Q3) Can concept-based guidance help mitigate RL caveats of SCoBot agents, such as policy misalignments?
@@
Figure 4: SCoBots allow to follow their decision process, because of their object centricity and interpretable concepts. The decision processes of SCoBots (extracted from the decision trees) with the associated states from a guided SCoBot on Skiing (left), and from normal SCoBots on Pong (middle) and Kangaroo (right). For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$. This agent selects the correct action for the right reasons.
Experimental setup: We evaluate SCoBots on $9$ Atari games (cf. Fig. 3 (left) for complete list) from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf. App. A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game (i.e. still able to return the ball).
For results based on achieved game scores we provide human normalized scores (following eq. 6) that are averaged over $3$ seeds for each agent configuration. We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following. Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in (Schulman et al., 2017), however we obtain similar results when evaluating on the old v4 versions (cf. Tab. 1). to prevent overfitting. All agents are trained for $20\text{M}$ frames under the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 implementation (Raffin et al., 2021) and its default hyperparameters (cf. Tab. 2 in App. A.5).
@@
We focus our SCoBot evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g. (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)). We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a). Specifically, in our evaluations, the set of object properties consists of object class (e.g. enemy paddle, ball etc.), $(x^{t},y^{t})$ coordinates, coordinates at the previous position $(x^{t-1},y^{t-1})$, height and width, the most common RGB values (i.e. the most representative color of an object), and object orientation. The specific set of functions, $\mathcal{F}$, that are used to extract object relations is composed of: euclidean distance, directed distances (on $x$ and $y$ axes), speed, velocity, plan intersection, the center (of the shortest line between two objects), and color name (cf. App. A.5.1 for implementation details).
To distill SCoBots’ learned policies, we use a rule-based model extraction method, similar to (Zarlenga et al., 2021). For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf. Tab. 4). Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below. More details about the experimental setup and the hyperparameters are provided in App. A.5.
SCoBots learn competitive policies (Q1).
We present human-normalized (HN) scores of both SCoBot and deep agents trained on each investigated game individually in Fig. 3 (left). Numerical values are provided in Tab. 1 (cf. App. A.7). We observe that SCoBot agents perform at least on par with deep agents on all games, even slightly outperform these on $5$ out of $9$ (namely, Asterix, Boxing, Kangaroo, Seaquest and Tennis). Our results suggest that SCoBots provide competitive performances on every tested game despite the constraints of multiple bottlenecks within their architectures. Overall, our experimental evaluations show that RL policies based on interpretable concepts extraction decoupled from the action selection process can, in principle, lead to competitive agents.
Inspectable SCoBots’ to detect misalignments (Q2).
@@
The main target of developing ScoBots is to obtain competitive, yet transparent agents that allow for human users to identify their underlying reasoning process. Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allows to pinpoint not just the object properties, but importantly the relational concepts being used for the final action decision. This is exemplified in Fig. 4 on Skiing, Pong and Kangaroo (cf. App. A.4 for explanations of SCoBots on the remaining games). Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of each game. For example, for the specific game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels). Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
@@
A more striking display of the benefits of the inherent explanations of SCoBot agents is depicted by the Pong agent in Fig 4. Specifically, the provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis). In fact, this suggests that the agent is largely ignoring the ball’s position. Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay). Thus, the vertical positions of these two objects are highly correlated (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf. App. A.7). Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.
@@
To validate these findings further, we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (referred to as NoEnemy cf. Fig. 1). We reevaluate the deep and SCoBot agents that were trained on the original Pong environment on NoEnemy and observe catastrophic performance drops in HN scores (cf. Fig. 3) at a level of random policies for both types of agents. Evaluations with different DQN agents lead to similar performance drops (cf. App. A.7). These drops are particularly striking as initial importance maps produced by PPO and DQN agents highlight all $3$ moving objects (i.e. the player, ball, and enemy) as relevant for their decisions (cf. Fig.1 top left and App. A.7) which is aligned with findings on importance maps of (Weitkamp et al., 2018). These maps suggest that the deep agents had based its decisions on right reasons. Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and granularity of such post-hoc explainability methods in the context of RL. Specifically, they highlight the importance of inherently explainable RL models on the level of concepts, as provided via SCoBots, to identify such potential issues. In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
@@
We here make use of the interactive properties of SCoBots to address several famous RL specific problems: goal misalignment, ill-defined rewards, difficult credit assignment and reward sparsity.
@@
Realigning SCoBots: For mitigating goal misalignment issues of the SCoBot agents trained on Pong, we can simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$). By doing so, the enemy cannot be used for the action selection process. This leads to guided SCoBots that are able to play Pong and its NoEnemy version (cf. guided SCoBot in Fig 3 (center)). Furthermore, this shows that playing Pong without observing the enemy is achievable, by simply returning vertical shots, difficult for the enemy to catch.
@@
Ill-defined reward: Defining a reward upfront that incentivizes agents to learn an expected behavior is a difficult problem. An example of ill-defined reward (i.e. a reward that will lead to an unaligned behavior if the agent maximizes the return) is present in Kangaroo. According to the documentation of the game333www.retrogames.cz/play_195-Atari2600.php, “the mother kangaroo on the bottom floor tries to reach the top floor where her joey is being held captive by some monkeys”. However, punching the monkey enemies gives a higher cumulative reward than climbing up to save the baby. RL agents thus tend to learn to kick monkeys on the bottom floor rather than reaching the joey (cf. Fig. 4). For revising such agents, we provide an additional reward signal, based on the distance from the mother kangaroo to the joey (detailed in App. A.8). As can be seen in Fig. 3 (right), this reward allows the guided SCoBots to achieve the originally intended goal by indeed completing the level. In contrast, deep agents and particularly unguided SCoBots achieve relatively high HN scores, but do not complete the levels.
@@
Difficult Credit Assignment Problem: The difficult credit assignment problem is the challenge of correctly attributing credit or blame to past actions when the agent’s actions have delayed or non-obvious effects on its overall performance. We illustrate this in the context of Skiing, where in standard configurations, agents receive at each step a negative reward that corresponds to the number of seconds elapsed since the last steps (i.e. varying between 3-3 and 7-7). This reward aims at punishing agents for going down the slope slowly. Additionally, the game keeps track of the number of flag pairs that the agent has correctly traversed (displayed at the top of the screen, cf. Fig. 4), and provide a large reward, proportional to this number, at the end of the episode. Associating this reward signal with the number of passed flags is extremely difficult, without prior knowledge on human skiing competitions. In the next evaluations, we provide SCoBots with another reward at each timestep, proportional to the agent’s progression to the flags’ position to incentivize the agent to pass in between them, as well as a signal rewarding for higher speed:
$\begin{split}R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}&-D(Player,o)^{t-1}\\[-8.61108pt] &+V(Player).y.\end{split}$ (5)
@@
These rewards are further detailed in App. A.8. With this, we obtain guided SCoBot agents that reach human scores, whereas the deep and unguided SCoBot agents perform worse than random (cf. Fig. 3 (left)).
Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf. App. A.8.3 for a detailled explanation).
@@
Overall, our experimental evaluations not only show that SCoBots are on par with deep agents in terms of performances, but that their concept-based and inspectable nature allows to identify and revise several important RL specific caveats. Importantly, in the course of our evaluations, we identified a previously unknown and critical misalignment of existing RL agents on the simple and iconic Pong environment, via the previously mentioned properties of SCoBots.
4 Limitations
In our evaluations, we made use of the integrated object extractor ($\omega$) of OCAtari which provides us with ground truth data. Such extractors can however also be optimized using supervised (Redmon et al., 2016) or self-supervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods. Furthermore, our set of relational functions, $\mathcal{F}$, is predefined and fixed. To eliminate the reliance on expert users, this set may alternatively be obtained using the supervision from large language models (Wu et al., 2023), via prior task-knowledge integration (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b). Other environments require additional information extraction processes. E.g. in MsPacman an agent must navigate a maze. Providing concept representations for such environments is an important step for future work. Finally, during our revision assessments, we presuppose that the user has adequate knowledge and positive intentions.